OpenStudy (anonymous):

can someone check my answer -2X(3y-4)-(xy+y)+3(x+y) =

Directrix (directrix):

>>can someone check my answer What did you get as an answer? @mdcaulfield

OpenStudy (anonymous):

did you add it up yet or is it already been answered

OpenStudy (anonymous):

Sorry - the answer I got is 9x+4y

Directrix (directrix):

@mdcaulfield -2 X (3y-4)-(xy+y)+3(x+y) = Are you thinking of the cap X as the variable x or as a symbol for multiplication in this problem?

Directrix (directrix):

>> the answer I got is 9x+4y Not what I got but I used -2x(3y-4)-(xy+y)+3(x+y) = and not this: -2X(3y-4)-(xy+y)+3(x+y) =